OpenStudy (anonymous):

How do I find the legs of an Isosceles Right Triangle when the given is the Hypotenuse and it is 5√6?

OpenStudy (raden):

x sqrt(2) = 5 sqrt(6) solve for x x is the leg of an isosceles right triangle

OpenStudy (anonymous):

or by pythag ... a = b soo a^2 + a^2 = c^2 2 a^2 = c^2 c = sqrt(2) a, solve for a

OpenStudy (anonymous):

ok so would it be √2 divided by 2 or just 2√2=2.8?